# LightCode: Light Analytical and Neural Codes for Channels with Feedback

### v2

@@ IV Proposed analytical coding scheme: PowerBlast
In this section, we propose PowerBlast, a hybrid 2-phase scheme, to iteratively refine the LMMSE estimate of the PAM symbol at the receiver and shift to a discrete symbol scheme that takes advantage of the sparsity of the error in the estimate of the PAM index in the final round. Through theoretical analysis and empirical results, we demonstrate that the performance of PowerBlast is better than both SK and GN, in several regimes of interest.
For a rate $\nicefrac{{K}}{{D}}$ code, PowerBlast begins by mapping $K$ bits of information to a $2^{K}$ PAM constellation and transmits the symbols on the forward AWGN channel. The receiver performs an LMMSE estimate and sends the estimate as feedback to the transmitter through the noiseless feedback channel. This continues for the first $D-1$ rounds. In the final round, PowerBlast uses a discrete symbol strategy to send the error in the PAM index of the estimate. The complete algorithm is described in Alg. 2.
Round 1: Tx: Power normalization: $x_{1}=\sqrt{P}\Theta$;
Forward channel: $y_{1}=x_{1}+n_{1}$;
Rx: LMMSE estimate of transmit symbol $\hat{\Theta}_{1}=\frac{y_{1}}{1+\sigma_{ff}^{2}/P}$ ;
/* Tx sends the error in estimate $\hat{\Theta}_{1}-\Theta$ over the next $D-2$ rounds */
while $12\leq i\leq D-$ 1 do
       Tx: Compute the error in estimate of previous round ${\epsilon}_{i-1}=\hat{\Theta}_{i-1}-\Theta$ ;
       Power normalization: $x_{i}=\frac{\sqrt{P}}{\sigma_{i-1}}{\epsilon}_{i-1}$, where $\sigma^{2}_{i-1}=\mathbb{E}[\epsilon^{2}_{i-1}]$;
       Forward channel: $y_{i}=x_{i}+n_{i}$;
       Rx: LMMSE estimate of transmit symbol $\hat{\epsilon}_{i}=\frac{\sqrt{P}\sigma_{i-1}}{P+\sigma_{ff}^{2}}y_{i}$ ;
       Update the estimate of $\Theta$ as: $\hat{\Theta}_{i}=\hat{\Theta}_{i-1}-\hat{\epsilon}_{i-1}$
end while
Decoding after round $D-1$: Map $\hat{\Theta}_{D-1}=\frac{\hat{x_{1}}}{\sqrt{P_{1}}\eta}$ to the closest symbol in the $2^{K}$ PAM constellation.
/* High-SNR scheme for final round */
Round D: Tx: Compute the difference in PAM indices $U=\hat{m}-m$;
Power normalization: $x_{D}=\frac{\sqrt{P}U}{\sigma_{ff}}$; where $U=\hat{m}-m$;
Transmit: $y_{D}=x_{D}+n_{D}$;
Final Decoding: Use ML decoder to detect $\hat{U}$ and detect original PAM signal using $\hat{U}$.
@@
We note that PowerBlast can be seen as a combination of SK and GN. The key difference between the SK scheme and the initial rounds of the GN scheme lies in the choice of information transmitted in the second round. In SK, the difference between $\Theta$ and the LMMSE estimate $\hat{\Theta}$ is transmitted, whereas in GN, the noise in the channel realization from round 1, $\mathbf{n_{1}}$, is transmitted. Hence, PowerBlast can be interpreted as $D-1$ rounds of SK scheme followed by GN scheme for the final round, resulting in a non-negligible improvement in performance. The error analysis for the PowerBlast scheme is provided below.
@@
We begin by computing the probability of error for phase 1, which consists of $D-1$ rounds of SK scheme. For ease of analysis, we follow the same assumption of  [5], where the estimator after round one is assumed to be the minimum variance unbiased estimator instead of LMMSE implying,
@@
Further, based on Alg. 1, we can view the effective channel corresponding to rounds 1 to $D-1$ as one round with effective SNR $S.1+S)^{D-2}$, where $S=\nicefrac{{P}}{{\sigma_{ff}^{2}}}$ is SNR for the forward channel in linear scale. Further, it is straightforward to show that the probability of error incurred in transmitting a symbol from unit power normalized $2^{K}$ PAM, using an optimal detector, is
@@
We now proceed to compute the probability of error for phase two, which is transmitting the difference in integer index over an AWGN channel. The key assumption for the high-SNR region is that the noise variance is small enough (compared to the effective forward SNR) that any errors beyond decoding to adjacent PAM symbols i.e., errors beyond 1 index difference are essentially negligible [4]. Hence, the problem can be rephrased as finding the probability of error to communicate a symbol from the constellation $\{-1,0,1\}$, with probability distribution $\{\nicefrac{{p_{SK}}}{{2}},1-p_{SK},\nicefrac{{p_{SK}}}{{2}}\}$ using MAP decoder. This can be computed as
@@
PowerBlast vs. SK and GN schemes. The performance comparison between the error expressions for a general SNR and rate is not straightforward (as they are represented in terms of Q functions). Instead, we limit our comparison to the canonical settings considered in the recent literature on channels with feedback (e.g., [26]), for rates $\nicefrac{{3}}{{6}}$ and $\nicefrac{{3}}{{9}}$ and plot the BLER performance in Fig. LABEL:fig:bler_gn_vs_pb; we observe significant gains of PowerBlast compared to both SK and GN in terms of the BLER performance.
@@
Figure 2: By combining SK and discrete-symbol strategy of GN scheme, PowerBlast noticeably improves the BLER performance upon both SK and GN schemes.
@@ V-A RNN families
deep-learning-based algorithms are capable of modeling complex input-output relations. One of the key challenges in designing codes for channels with feedback is accurately computing the next transmission that minimizes the probability of decoding error, conditioned on the feedback from previous rounds. Existing classical coding schemes employ linear estimators at the receivers and model the dependencies in a linear fashion at the transmitter, which is sub-optimal. Instead, the sequential nature of the feedback from previous can be better leveraged by using deep-learning architectures such as RNNs that are tailored for processing sequential data. Deepcode [21] demonstrated this advantage in modeling the dependencies across bits and rounds using bi-directional gated recurrent units (GRUs). Several follow-up works investigated the use of other RNN-based architectures such as long short-term memory (LSTMs) [22, 23] and, more recently, Robust Coding [24] was proposed, which combined attention mechanism with bi-directional GRUs to optimize the symbol-by-symbol code design for noisy feedback channels across the rounds. Despite the promising performance, one disadvantage of using RNN-based architecture is the necessity to store the hidden states of the model, which are typically of much higher dimensions than the inputs and demand a lot of memory. Additionally, the sequential and iterative nature of encoding and decoding in RNNs can result in significant latency in the system.
@@ V-B Transformer families
In another line of work, self-attention-based transformer architectures have been explored for designing neural feedback codes. AttentionCode [25] introduced the idea of replacing RNN-based architecture with pure attention-based models, resulting in better alignment between a symbol and the corresponding feedback. More recently, GBAF [26] introduced the idea of performing block coding across the codeblock, in addition to coding across the rounds, resulting in orders of magnitude improvement in performance at extremely low SNRs.
As illustrated in Fig. 3, the encoding of GBAF is performed across the rounds by causally concatenating the message and feedback symbols from previous rounds and using a series of feature extractors and multi-layer perception (MLP) modules. Additionally, positional encoding (PE) and a self-attention-based transformer encoder layer are deployed to encourage the mixing of information across the symbols within a codeblock, leading to block coding. More specifically, a block of $L$ bits is divided into $l$ sub-blocks of $K$ bits each and first encoded independently using a feature extractor. Next, PE and transformer encoder modules perform cross-symbol coding across the $l$ symbols. Finally, an MLP module is used to encode each symbol. A similar architecture is used at the decoder.
@@ V-C Important open problems
While providing impressive performance in reliability, transformer architecture is computationally expensive. Further, it is well known that the transformer architecture does not scale well to larger sequences. The self-attention mechanism imposes a compute complexity of $O(n^{2})$ during training and $O(n)$ during inference, even after using KV cache, with respect to input length $n$. Moreover, as the blocklength scales, a memory complexity of $O(n^{2})$ prohibits training the algorithm from using a large batch size, which is crucial for attaining a good performance for any deep-learning models.
@@ VI Proposed neural coding scheme: LightCode
Our goal is to design lightweight neural codes without the need for block coding i.e., we limit ourselves to symbol-by-symbol schemes suitable for both noiseless and noisy feedback settings. To this end, we design a lightweight deep-learning-based scheme with 10×10$\times$10 × fewer parameters compared to existing schemes. Surprisingly, this low-complex solution achieves a performance superior to current state-of-the-art deep-learning-based block-coding schemes. We now present the architecture and training choices that were crucial in achieving this new state-of-the-art BLER performance.
@@ VI-A LightCode: Architecture
Our architecture. We split the design of the encoder-decoder architecture into two parts: the feature extractor and the MLP module, as illustrated in Fig. 3 (right). The choice of the feature extractor plays a crucial role in determining the downstream performance. A higher complexity feature extractor can provide a better performance but might not be desirable for practical applications. For LightCode, after experimenting with various choices, we found that the design illustrated in Fig. 4 gave the optimal trade-off in complexity vs BLER performance. Specifically, we consider a hidden dimension of $d=32$ and 3 hidden layers. At the final layer, we use a skip connection to aggregate the outputs of layer 1 and layer 3 to project down to a latent dimension of $d=16$. The motivation for this architecture comes from our prior work on the design of deep-learning-based non-linear polar codes, DeepPolar [18], where skip connections have proven to be crucial to ensure good performance. Finally, the output of the feature extractor will be passed to an MLP module to perform encoding or decoding. For the encoder, we choose a $2$ layer MLP, and for the decoder, we choose a $1$ layer MLP to transform the features into encoded representations or decoder outputs, respectively.
@@
Our architecture vs. transformers. While transformer models overcome the issue of storing hidden states by using positional encoding and self-attention, they also come with great computational complexity. Moreover, recent transformer-based feedback coding schemes proposed block coding using self-attention, which requires $O(n^{2})$ memory with respect to codeblock length. Hence, a feed-forward architecture to perform a symbol-by-symbol coding scheme is more practical.
By carefully designing the architecture suitable for a symbol-by-symbol scheme, we achieve a lightweight design with more than 10×10$\times$10 × reduction in the number of parameters compared to the RNN family of schemes such as Robust Coding and transformer family of block coding schemes such as GBAF. Further, we show in Section VIII-C that this reduction in complexity provides up to 171×171$\times$171 × higher decoding throughput compared to Robust Coding and up to 10×10$\times$10 × higher throughput compared to GBAF. We now describe the encoding and decoding procedure in detail below.
@@
$x_{i}=\alpha_{i}\phi(\mathbf{u},\tilde{y}_{1},\tilde{y}_{2},\dots,\tilde{y}_{i-1}),$ (18)
where $\tilde{y}_{i}=x_{i}+n_{i}+\tilde{n}_{i}$ is the feedback from the previous round and $\alpha_{i}$ is the scaling factor to ensure sum power constraint across the rounds $\sum_{i=1}^{D}\alpha_{i}^{2}=D$.
@@
Choice of feature vectors. Empirically, we observed that the choice of input to the feature extractor has a noticeable effect on the BLER performance based on the feedback channel. For noiseless feedback, the input features $(\mathbf{u},y_{1},y_{2},\dots,y_{i})$ worked the best. Further, as will be evident from discussions in Section VII-C, the input features $(\mathbf{u},x_{1},x_{2},\dots,x_{i},n_{1}+\tilde{n}_{1},n_{2}+\tilde{n}_{2},\dots n_{i}+\tilde{n}_{i})$ gave the best performance for noisy feedback. Apart from this difference, the rest of the architecture remains the same for noiseless and noisy feedback scenarios.
Figure 4: Feature extractor design for LightCode.
Support for multiple rates. One of the limitations of existing symbol-by-symbol coding schemes is the inability to serve a variety of rates i.e., only rates of the form $\nicefrac{{1}}{{D}},D\in\mathbb{Z^{+}}$ are supported. In order to overcome this, instead of processing 1 bit at a time, LightCode encodes a block of $K$ bits into 1 symbol and communicates this symbol to the receiver over $D$ rounds. By independently varying $K$ and $D$ any rate of the form $\nicefrac{{K}}{{D}}$ where, $K,D\in\mathbb{Z^{+}}$ can be supported, making the scheme more flexible, while simultaneously reducing the overall latency of the communication by a factor of $K$. The results for multiple rates for a block length of $K=3$ are discussed in Table II.

### v3

@@ IV Proposed analytical coding scheme: PowerBlast
In this section, we propose PowerBlast, a hybrid 2-phase scheme, to iteratively refine the LMMSE estimate of the PAM symbol at the receiver and shift to a discrete symbol scheme that takes advantage of the sparsity of the error in the estimate of the PAM index in the final round. Through theoretical analysis and empirical results, we demonstrate that the performance of PowerBlast is better than both SK and GN in several regimes of interest.
For a rate $\nicefrac{{K}}{{D}}$ code, PowerBlast begins by mapping $K$ bits of information to a $2^{K}$ PAM constellation and transmits the symbols on the forward AWGN channel. The receiver performs an LMMSE estimate and sends the estimate as feedback to the transmitter through the noiseless feedback channel. This continues for the first $D-1$ rounds. In the final round, PowerBlast uses a discrete symbol strategy to send the error in the PAM index of the estimate. The complete algorithm is described in Algorithm 2.
1
Input: Message symbol $\Theta$, number of rounds $D$, forward noise variance $\sigma_{ff}^{2}$
2
3 Round 1: Tx: Power normalization: $x_{1}=\sqrt{P}\Theta$;
4 Forward channel: $y_{1}=x_{1}+n_{1}$;
5 Rx: LMMSE estimate of transmit symbol $\hat{\Theta}_{1}=\frac{\sqrt{P}y_{1}}{P+\sigma_{ff}^{2}};$
6 /* Tx communicates the error in estimate $\hat{\Theta}_{1}-\Theta$ over the next $D-2$ rounds */
7
8while $12\leq i\leq D-$ 1 do
9       Tx: Compute the error in estimate of previous round ${\epsilon}_{i-1}=\hat{\Theta}_{i-1}-\Theta$ ;
10       Power normalization: $x_{i}=\frac{\sqrt{P}}{\sigma_{i-1}}{\epsilon}_{i-1}$, $\sigma^{2}_{i-1}=\mathbb{E}[\epsilon^{2}_{i-1}]$;
11       Forward channel: $y_{i}=x_{i}+n_{i}$;
12       Rx: LMMSE estimate of transmit symbol $\hat{\epsilon}_{i}=\frac{\sqrt{P}\sigma_{i-1}}{P+\sigma_{ff}^{2}}y_{i}$ ;
13       Update the estimate of $\Theta$ as: $\hat{\Theta}_{i}=\hat{\Theta}_{i-1}-\hat{\epsilon}_{i-1}$
14
15 Decoding after round $D-1$: Map $\hat{\Theta}_{D-1}=\frac{\hat{x_{1}}}{\sqrt{P_{1}}\eta}$ to the closest symbol in the $2^{K}$ PAM constellation.
16 /* High-SNR scheme for final round */
17
18Round D: Tx: Compute the difference in PAM indices $U=\hat{M}-M$, where $\hat{M}$ and $M$ correspond to the integer index from PAM constellation for $\hat{\Theta}$ and $\Theta$ respectively;
19 Power normalization: $x_{D}=\frac{\sqrt{P}U}{\sigma_{ff}}$; Transmit: $y_{D}=x_{D}+n_{D}$;
20 Final Decoding: Use MAP decoder to detect $\hat{U}$ and detect original PAM signal using $\hat{U}$.
@@ 19 Power normalization: $x_{D}=\frac{\sqrt{P}U}{\sigma_{ff}}$; Transmit: $y_{D}=x_{D}+n_{D}$;
We note that PowerBlast can be seen as a combination of SK and GN. Rate $\nicefrac{{K}}{{D}}$ PowerBlast can be interpreted as $D-1$ rounds of rate $\nicefrac{{K}}{{D}}$ SK scheme with LMMSE estimate, followed by the discrete variant of GN scheme for the final round, resulting in a non-negligible improvement in performance in certain regions of SNR. The error analysis for the PowerBlast scheme is provided below.
@@
We begin by computing the probability of error for phase 1, which consists of $D-1$ rounds of the SK scheme. For ease of analysis, we follow the same assumption of  [5], where the estimator after round one is assumed to be MVUE instead of LMMSE, implying,
@@
Further, based on Algorithm 1 , we can view the effective channel corresponding to rounds 1 to $D-1$ as one round with effective SNR $S.(1+S)^{D-2}$, $S=\nicefrac{{P}}{{\sigma_{ff}^{2}}}$ is SNR for the forward channel in linear scale. It is straightforward to show that the probability of error incurred in transmitting a symbol from unit power normalized $2^{K}$ PAM, using an optimal detector, is
@@
We now proceed to compute the probability of error for phase two, which is transmitting the difference in integer (message) index over an AWGN channel. The key assumption for the high-SNR region is that the noise variance is small enough (compared to the effective forward SNR) that any errors beyond decoding to adjacent PAM symbols i.e., errors beyond 1 index difference are essentially negligible [4]. Hence, the problem can be rephrased as finding the probability of error to communicate a symbol from the constellation $\{-1,0,1\}$, with probability distribution $\{\nicefrac{{p_{SK}}}{{2}},1-p_{SK},\nicefrac{{p_{SK}}}{{2}}\}$ using MAP decoder. This can be computed as
@@
PowerBlast vs. SK and GN schemes. The key difference between PowerBlast and SK is in the last round of communication, where PowerBlast shifts to a discrete symbol scheme, as seen in line 13 of Algorithm 2, by taking advantage of the high effective SNR over the initial $D-1$ rounds i.e., the (effective) noise variance is much smaller than the distance between adjacent PAM symbols. Hence, it is beneficial to now transmit the error in the PAM constellation index, which lies in $\{-1,0,1\}$ with a very high probability, resulting in a much lower probability of decoding error.
Further, compared to GN, the key difference in PowerBlast is the choice of information transmitted after round 1. GN communicates and refines the noise from round 1, $n_{1}$, starting round 2, as seen in line 4 of Algorithm 4. In contrast, PowerBlast transmits and refines the error in LMMSE estimate of $\Theta$, as seen in line 4 of Algorithm 2. It is seen from [5] that the effective SNR (in linear scale) after $D$ rounds is given by $S(1+S)^{D-1}$ for GN and $(1+S)^{D}-1$ for PowerBlast. While the difference in effective SNR becomes negligible for large $D$, it can be significant when operating for a finite number of rounds, resulting in the superior performance of PowerBlast.
The performance comparison between the error expressions for a general SNR and rate is not straightforward (as they are represented in terms of Q functions). Instead, we limit our comparison to the canonical settings considered in the recent literature on channels with feedback (e.g., [26]), for rates $\nicefrac{{3}}{{6}}$ and $\nicefrac{{3}}{{9}}$ and plot the BLER performance in Fig. 2; we observe significant gains of PowerBlast compared to both SK and GN in terms of the BLER performance.
@@
Figure 2: By combining the SK and discrete-symbol strategy of the GN scheme, PowerBlast noticeably improves the BLER performance upon both SK and GN schemes.
@@ V-A RNN families
Deep-learning-based algorithms are capable of modeling complex input-output relations. One of the critical challenges in designing codes for channels with feedback is accurately computing the subsequent transmission that minimizes the probability of decoding error, conditioned on the feedback from previous rounds. Existing classical coding schemes employ linear estimators at the receivers and model the dependencies in a linear fashion at the transmitter, which is sub-optimal. Instead, the sequential nature of the feedback from previous can be better leveraged by using deep-learning architectures such as RNNs tailored for processing sequential data. Deepcode [21] demonstrated this advantage in modeling the dependencies across bits and rounds using bi-directional gated recurrent units (GRUs). Several follow-up works investigated the use of other RNN-based architectures such as long short-term memory (LSTMs) [22, 23] and, more recently, Robust Coding [24] was proposed, which combined attention mechanism with bi-directional GRUs to optimize the symbol-by-symbol code design for noisy feedback channels across the rounds. Despite the promising performance, one disadvantage of using RNN-based architecture is the necessity to store the hidden states of the model, which are typically of much higher dimensions than the inputs and demand a lot of memory. Additionally, the sequential and iterative nature of encoding and decoding in RNNs can result in significant latency in the system.
@@ V-B Transformer families
In another line of work, self-attention-based transformer architectures have been explored for designing neural feedback codes. AttentionCode [25] introduced the idea of replacing RNN-based architecture with pure attention-based models, resulting in better alignment between a symbol and the corresponding feedback. In other words, AttentionCode can be viewed as Deepcode with transformer architecture. By leveraging the attention mechanism, AttentionCode creates temporal correlations at the transmitter for encoding and exploits these temporal correlations at the receiver for decoding. Further, the inputs to the encoder and decoder transformers are restructured to align each bit with the corresponding noise from multiple rounds as a single column, processed by the self-attention mechanism. More recently, GBAF [26] introduced the idea of performing block coding across the codeblock, in addition to temporal coding across the rounds, resulting in orders of magnitude improvement in performance at extremely low SNRs, explained in detail below.
As illustrated in Fig. 3, the encoding of GBAF is performed across the rounds by causally concatenating the message and feedback symbols from previous rounds and using a series of feature extractors and multi-layer perception (MLP) modules. Additionally, positional encoding (PE) and a self-attention-based transformer encoder layer are deployed to encourage the mixing of information across the symbols within a codeblock, leading to block coding. More specifically, a block of $L$ bits is divided into $l$ sub-blocks of $K$ bits each and first encoded independently using a feature extractor. Next, PE and transformer encoder modules perform cross-symbol coding across the $l$ symbols. Finally, an MLP module is used to encode each symbol. A similar architecture is used at the decoder, but the output dimensions are adjusted accordingly.
@@ V-C Important open problems
While providing impressive performance and reliability, transformer architecture is computationally expensive. Further, it is well known that the transformer architecture does not scale well to larger sequences. The self-attention mechanism imposes a compute complexity of $O(n^{2})$ during training and $O(n)$ during inference, even after using the KV cache, with respect to input length $n$. Moreover, as the blocklength scales, a memory complexity of $O(n^{2})$ prohibits training the algorithm from using a large batch size, which is crucial for attaining a good performance for any deep-learning models.
@@ VI Proposed neural coding scheme: LightCode
Our goal is to design lightweight neural codes without the need for block coding i.e., we limit ourselves to symbol-by-symbol schemes suitable for both noiseless and noisy feedback settings. To this end, we design a lightweight deep-learning-based scheme with 10×10$\times$10 × fewer parameters compared to existing schemes. Surprisingly, this low-complex solution achieves a performance superior to current state-of-the-art deep-learning-based block-coding schemes. We now present the architecture and training choices crucial to achieving this new state-of-the-art BLER performance.
@@ VI-A LightCode: Architecture
Our architecture. We split the design of the encoder-decoder architecture into two parts: the feature extractor and the multi-layer perception (MLP) module, as illustrated in Fig. 3 (right). The choice of the feature extractor plays a crucial role in determining the downstream performance. A higher complexity feature extractor can perform better but might not be desirable for practical applications. For LightCode, after experimenting with various choices, we found that the design illustrated in Fig. 4 gave the optimal trade-off in complexity vs. BLER performance. Further, the output of the feature extractor will be passed to an MLP module. For the encoder, we choose a $1$ layer MLP to project the features to a $1$ dimensional output. For the decoder, we choose a $2$ layer MLP to transform the features into an output of dimension $2^{K}$. The full architecture for encoder and decoder, including the MLP are provided in Appendix B, Fig. 14.
@@
Our architecture vs. transformers. While transformer models overcome the issue of storing hidden states by using positional encoding and self-attention, they also come with great computational complexity. Moreover, recent transformer-based feedback schemes proposed block coding using self-attention, which requires $O(n^{2})$ memory and compute with respect to codeblock length. But as evident from results Section VII, block coding does not seem to provide noticeable gains in the setting under consideration. Hence, we propose a symbol-by-symbol coding scheme that uses a significantly simpler feed-forward architecture.
Complexity. By eliminating the need for block coding and using a short code length, LightCode avoids the high-complexity transformer module used in GBAF architecture and instead uses a simple feed-forward network. By carefully designing the architecture suitable for a symbol-by-symbol scheme, we achieve a lightweight design with more than 10×10$\times$10 × reduction in the number of parameters compared to the RNN family of schemes such as Robust Coding and transformer family of block coding schemes such as GBAF. Further, we show in Section VIII-C that this reduction in complexity provides up to 171×171$\times$171 × higher decoding throughput compared to Robust Coding and up to 10×10$\times$10 × higher throughput compared to GBAF.
Feature extractor. Fig. 4 depicts the architecture of the feature extractor, which is the backbone for both encoder and decoder models. Compared to the feature extractor used in GBAF, we introduce two changes to improve the performance and reduce the complexity. The first is to add a skip connection, which preserves the prior from input to the feature extractor better, similar to the design of DeepPolar [18]. Further, the hidden dimension is reduced from $64$ to $32$, decreasing the number of parameters.
Specifically, the input to the feature extractor is processed through a sequence of three linear layers, each with a hidden dimension of $32$, and rectified linear unit (ReLU) activation functions are applied between the first two layers. In parallel, we add a skip connection from output of first layer to input of final layer, which reverses the sign of the representation to introduce variance in the information. These outputs are concatenated and passed through the final layer, which generates a 16-dimensional feature representation.
Training. Short code length and lightweight architecture allow for training of LightCode with an extremely large batch size of $10^{5}$, which is more than 10×10$\times$10 × larger than the maximum batch size suitable for GBAF codes. We refer to Section VI-B for a detailed discussion of the training details and Section VIII-A for an ablation study on the role of batch size.
We now describe the detailed encoding and decoding procedure below.
@@
$x_{i}=\alpha_{i}\phi(\mathbf{u},\tilde{y}_{1},\tilde{y}_{2},\dots,\tilde{y}_{i-1},0,\dots,0),$ (18)
where $\tilde{y}_{i}=x_{i}+n_{i}+\tilde{n}_{i}$ is the feedback from the previous round and $\alpha_{i}$ is the scaling factor to ensure sum power constraint across the rounds $\sum_{i=1}^{D}\alpha_{i}^{2}=D$. Here, we note that in order to keep the size of the input to the encoder constant, the input is padded with zeros where necessary.
@@
Choice of feature vectors. Empirically, we observed that the choice of inputs to the feature extractor has a noticeable effect on the BLER performance based on the feedback channel. For noiseless feedback, the input features $(\mathbf{u},y_{1},y_{2},\dots,y_{i})$ worked the best. Further, as will be evident from discussions in Section VII-C, the input features $(\mathbf{u},x_{1},x_{2},\dots,x_{i},n_{1}+\tilde{n}_{1},n_{2}+\tilde{n}_{2},\dots n_{i}+\tilde{n}_{i})$ gave the best performance for noisy feedback. Apart from this difference in features, the rest of the architecture remains the same for noiseless and noisy feedback scenarios.
Figure 4: Feature extractor design for LightCode for a rate $\nicefrac{{3}}{{9}}$ code.
Support for multiple rates. One of the limitations of existing symbol-by-symbol coding schemes is the inability to serve a variety of rates i.e., only rates of the form $\nicefrac{{1}}{{D}},D\in\mathbb{Z^{+}}$ are supported. In order to overcome this, instead of processing 1 bit at a time, LightCode encodes a block of $K$ bits into $1$ symbol and communicates this symbol to the receiver over $D$ rounds. By independently varying $K$ and $D$ any rate of the form $\nicefrac{{K}}{{D}}$ where, $K,D\in\mathbb{Z^{+}}$ can be supported, making the scheme more flexible, while simultaneously reducing the overall latency of the communication by a factor of $K$. The results for multiple rates for a block length of $K=3$ are discussed in Table III.